A stone is dropped into a well 19.6m deep and the impact pf sound is heard after 2.056 seconds. Find the velocity of sound in air


1
Expert's answer
2021-03-26T11:36:11-0400

"t=t_1+t_2"


"h=gt_1^2/2\to t_1=\sqrt{2h/g}=\sqrt{2\cdot 19.6/9.81}\approx1.999\ (s)"


"t_2=h\/v"


"2.056=1.999+19.6/v\to v=19.6/0.057\approx344\ (m/s)" . Answer